OpenStudy (anonymous):

I need help with writing a parallel and perpendicular line equation for y=1/4x-2(8,-1) please

OpenStudy (anonymous):

think that in order for being parallel the slope has to be the same, if you want a perpendicular the new equations slope should be 1/m , were m is your slope.

OpenStudy (anonymous):

I need to simplify my equation into slope-intercept form

OpenStudy (anonymous):

the regular intercept slopt format is y = mx + b, therefore you should have something like this for your equation y = x/4 - 2 . im not sure why you give me that point thoug ( 8, -1)

OpenStudy (anonymous):

This point was already added to the equation when it was given to me

OpenStudy (anonymous):

the paralell equation is pretty easy, just pick any b since that would be the origin point. lets say y = x/4 + 9. If you graph that in www.wolframalpha.com and the one you gave me you can notice that both are parallel..

OpenStudy (anonymous):

We have to discuss the necessary steps to carry out the activity and describe what each line looks like in relation to the orginal given line.

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

yw, for the perpendicular you must notice that the slope is 1/4. Since the slope is 1/4 if you multuply that by 1/4 again you would get 4. the following equations slope should be 4. Taking the same original point the perpendicular equation would be y = 4x - 2. Graph it :)

OpenStudy (anonymous):

What about y=-4x-5; (0,-1)

OpenStudy (anonymous):

same thing, m = -4 , therefore the perpendicular is 1/m , in which case -1/4. parallel keeps the m and the b changes, just give it instead of -5 a 2.

OpenStudy (anonymous):

Thanks